OpenStudy (anonymous):

Evaluate (x + y)^0 for x = 4 and y = 3. a) 0 b) 1 c) 2 d)7

OpenStudy (anonymous):

remember order of operations: PEMDAS (Please Excuse My Dear Aunt Sally): Parentheses, Exponents, Multiplication/Division, Addition/Subtraction. So the first thing to do is what is inside the parentheses: (x+y) = (4+3) = 7. Next we do the exponent part: 7^0 = 1, because anything raised to the power 0 is 1.